OpenStudy (anonymous):

Write an equation of a line in point-slope form that has a slope of -2 and passes through (5, -1).

OpenStudy (anonymous):

\[y-(-1)=-2(x-5)\]

OpenStudy (anonymous):

If you need it explained I can do that :)

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

Happily most people just want answers!!! Hold on just one second

OpenStudy (anonymous):

Alright thank you

OpenStudy (anonymous):

Point-slope form looks like: \[(y-y _{1})=m(x-x _{1})\] where m=slope and \[(x _{1},y _{1})\] is the point given. So all you need to do is plug in the info they gave you. \[y-(-1)=-2(x-5)=y+1=-2(x-5)\]

OpenStudy (anonymous):

Oh im so slow. But thank you for explaining it for me

OpenStudy (anonymous):

No problem! I had to look up which one was point-slope form because I haven't seen it in a while.